# The physical and chemical processes in protoplanetary disks: constraints on the composition of comets

### v3

@@ 4.2 Estimation of volatile composition in solids
Observations of fresh sublimates are the most straightforward to reveal the ice composition, and are possible during luminosity outbursts. The outbursts, which are caused by a sudden increase of mass accretion from the disk to the central star, are observed in Class I or the transition phase of Class I to Class II (§2.1). While the water snowline is typically located at a few au in Class II disks, it extends further away upon outburst. Around V883 Ori the dust continuum observation suggests that the water snowline is located at au (Cieza et al., 2016). V883 Ori is a FU Ori type star, for which the duration of the outburst phase is years. Since it is shorter than typical timescales of the gas-phase chemistry, we can observe fresh sublimates. Lee et al. (2019) detected abundant COMs (e.g. CHCHO and CHOCHO); their abundances relative to CHOH are comparable to those in comets and hot corinos (see the chapter by Bergin et al.). A line of HDO is also tentatively detected. When the burst ceases, the COMs and water will again freeze on grain surfaces.
As described in §3.1, observations of Class II disks outside the CO snowline indicate that carbon and oxygen are depleted in the warm molecular layer, most probably due to the sedimentation and radial drift of ice-coated dust grains. Observations of molecules such as CH suggest a high C/O ratio () in the gas phase, which in turn suggests that the solid phase is oxygen-rich, e.g. with abundant water ice. The molecular composition of ice could then be determined by the competition of the timescales of dust growth, settling, and molecular evolution in the upper layers of the disk. If the grain growth and settling are fast enough, the ices inherited from molecular clouds might be stored and survive in the midplane, which is cold and shielded from UV and X-ray radiation. Interestingly, recent observations suggest that the growth and sedimentation of dust grains and CO depletion proceed over the short timescale from Class 0/I to II (Bergner et al., 2020; Zhang et al., 2020; Harsono et al., 2021; Ohashi et al., 2023).
@@ 5 Brief summary and future directions
While ALMA has revealed the thermal emission of dust with sub-arcsecond resolution ( au), derivation of the mass distribution of dust grains depends on the opacity, which, in turn, depends on the size distribution of grains. The combination of the observations at various wavelengths, polarization, and theoretical studies on dust properties are crucial (§2). In order to probe the dust in the inner radius, which tends to be optically thick even in millimeter wavelength, low-frequency observations by Band 1 ALMA and ngVLA are essential. These low-frequency instruments are also suited for observations of major volatile molecules such as NH and CHOH. In recent years, various ring and linear molecules are newly detected in the line surveys of molecular cloud TMC-1 at Green Bank Telescope and Yabe 40 m telescope at GHz (Cernicharo et al., 2020; McGuire et al., 2020). While the detection of new molecules in disks could be limited by sensitivity, since large molecules suitable for low-frequency observations tend to be frozen in ice mantle, disks in outburst could be a good target.
@@
In shorter wavelengths, the James Webb Space Telescope (JWST) is now providing a comprehensive view of inner disk chemistry in protoplanetary disks around stars of a wide range of masses in the near future. Active programs include the JWST Disk Infrared Spectroscopic Chemistry Survey (JDISCS, Pontoppidan et al., 2023), and the Mid-infrared Disk Survey (MINDS, Kamp et al., 2023). These surveys promise to yield inventories of a wide range of bulk volatile species and their isotopologues within the snowline, including water, organics, and nitrogen-bearing species. While the JWST observations will be able to survey disks around stars of all masses, its medium-resolution spectrometer will not be able to spectrally resolve individual lines. However, a new generation of high-resolution spectrometers on ground-based facilities will continue to offer complementary spectroscopy at high resolution. In particular the new CRIRES+ instrument (Dorn et al., 2014) on the European Southern Observatory Very Large Telescope (ESO-VLT), and in the future, the METIS instrument (Brandl et al., 2010) of the European Extremely Large Telescope and MICHI (Packham et al., 2018) on the Thirty Meter Telescope will provide sensitive 3-12 m spectroscopy of protoplanetary disk chemistry at high spectral and spatial resolution. Finally, the US 2020 Decadal Survey recommended either a far-infrared or X-ray probe class misssion in the 2030s, and several concepts to trace the physics and chemistry of volatiles around the snowline using far-infrared spectroscopy are under consideration (e.g., PRIMA, Glenn et al., 2023).
@@ Acknowledgments.
Y.A. acknowledges support by NAOJ ALMA Scientific Research Grant Numbers 2019-13B, Grant-in-Aid for Scientific Research (S) 18H05222, and Grant-in-Aid for Transformative Research Areas (A) 20H05844 and 20H05847.

### v4

@@ 4.2 Estimation of volatile composition in solids
Observations of fresh sublimates are the most straightforward to reveal the ice composition, and are possible during luminosity outbursts. The outbursts, which are caused by a sudden increase of mass accretion from the disk to the central star, are observed in Class I or the transition phase of Class I to Class II (§2.1). While the water snowline is typically located at a few au in Class II disks, it extends further away upon outburst. Around V883 Ori the dust continuum observation suggests that the water snowline is located at au (Cieza et al., 2016). V883 Ori is a FU Ori type star, for which the duration of the outburst phase is years. Since it is shorter than typical timescales of the gas-phase chemistry, we can observe fresh sublimates. Various COMs (e.g. CH3CHO and CH3OCHO) are detected; their abundances relative to CH3OH are comparable to those in comets (Lee et al., 2019; Yamato et al., 2024b). HDO is recently detected as well (Tobin et al., 2023) (section 3.4.1).When the burst ceases, the COMs and water will again freeze on grain surfaces.
As described in §3.1, observations of Class II disks outside the CO snowline indicate that carbon and oxygen are depleted in the warm molecular layer, most probably due to the sedimentation and radial drift of ice-coated dust grains. Observations of molecules such as C2H suggest a high C/O ratio () in the gas phase, which in turn suggests that the solid phase is oxygen-rich, e.g. with abundant water ice. The molecular composition of ice could then be determined by the competition of the timescales of dust growth, settling, and molecular evolution in the upper layers of the disk. If the grain growth and settling are fast enough, the ices inherited from molecular clouds might be stored and survive in the midplane, which is cold and shielded from UV and X-ray radiation. Interestingly, recent observations suggest that the growth and sedimentation of dust grains and CO depletion proceed over the short timescale from Class 0/I to II (Bergner et al., 2020; Zhang et al., 2020; Harsono et al., 2021; Ohashi et al., 2023).
@@ 5 Brief summary and future directions
While ALMA has revealed the thermal emission of dust with sub-arcsecond resolution ( au), derivation of the mass distribution of dust grains depends on the opacity, which, in turn, depends on the size distribution of grains. The combination of the observations at various wavelengths, polarization, and theoretical studies on dust properties are crucial (§2). In order to probe the dust in the inner radius, which tends to be optically thick even in millimeter wavelength, low-frequency observations by Band 1 ALMA and ngVLA are essential. These low-frequency instruments are also suited for observations of major volatile molecules such as NH3 and CH3OH. In recent years, various ring and linear molecules are newly detected in the line surveys of molecular cloud TMC-1 at Green Bank Telescope and Yabe 40 m telescope at GHz (Cernicharo et al., 2020; McGuire et al., 2020). While the detection of new molecules in disks could be limited by sensitivity, since large molecules suitable for low-frequency observations tend to be frozen in ice mantle, disks in outburst could be a good target.
@@
In shorter wavelengths, the James Webb Space Telescope (JWST) is now providing a comprehensive view of inner disk chemistry in protoplanetary disks around stars of a wide range of masses. Active programs include the JWST Disk Infrared Spectroscopic Chemistry Survey (JDISCS, Pontoppidan et al., 2024), and the Mid-infrared Disk Survey (MINDS, Kamp et al., 2023). These surveys promise to yield inventories of a wide range of bulk volatile species and their isotopologues within the snowline, including water, organics, and nitrogen-bearing species. While the JWST observations will be able to survey disks around stars of all masses, its medium-resolution spectrometer will not be able to spectrally resolve individual lines. However, a new generation of high-resolution spectrometers on ground-based facilities will continue to offer complementary spectroscopy at high resolution. In particular the new CRIRES+ instrument (Dorn et al., 2014) on the European Southern Observatory Very Large Telescope (ESO-VLT), and in the future, the METIS instrument (Brandl et al., 2010) of the European Extremely Large Telescope and MICHI (Packham et al., 2018) on the Thirty Meter Telescope will provide sensitive 3-12 m spectroscopy of protoplanetary disk chemistry at high spectral and spatial resolution. Finally, the US 2020 Decadal Survey recommended either a far-infrared or X-ray probe class misssion in the 2030s, and several concepts to trace the physics and chemistry of volatiles around the snowline using far-infrared spectroscopy are under consideration (e.g., PRIMA, Glenn et al., 2023).
@@ Acknowledgments.
Y.A. acknowledges support by NAOJ ALMA Scientific Research Grant Numbers 2019-13B, Grant-in-Aid for Scientific Research (S) 18H05222, (B) 24K00674, and Grant-in-Aid for Transformative Research Areas (A) 20H05844 and 20H05847.